• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Roedelbronn

    Log entry time 08:09:15 on June19,2002

    Entry number 86401

    keyword=Owl shift summary

    Crew on shift: Michael Roedelbronn, Bodo Reitz, Ole Hansen

    00:00 We have beam at 70 uA.
    00:12 From spot++ we see that raster size is ~1.4 x 3 vice the desired 2 x 2. For consistency among measurements we stopped the run to wait until the raster is restored to what it has been for previous runs. Called MCC.
    00:15 Called again to check on raster size. May take a few minutes as they are doing passdown.
    00:30 Raster looks OK now. not exactly 2x2 but consistent with what we've had (currently ~1.9 x 2.3).
    01:00 MCC still able to deliver only 70 uA. We were turned over that they would try to raise it to 100 uA as beam stability would support. We have had several short trips, but otherwise the beam has been good at ~70uA.
    02:00 MCC was trying to bring beam up to 80uA but then beam tripped off for several minutes.
    02:15 MCC trying to restore beam at a higher current. ...at 90 uA now. We're not sure where we'll end up, so we'll keep the prescales and the current run going until it ends. Then we'll adjust prescales to 90 from 70 uA values which are currently in effect.
    02:22 ROC 14 crash. Reset from Hall A Slow Controls.
    02:27 ROC 14 reset successful. Resuming data-taking with new prescale values.
    02:46 MCC is bringing beam up to 100 uA. As before, we'll keep the run going and readjust prescalers for the 100 uA values for the following run.
    04:48 ROC 14 crash during run 1636. Reset from Hall A Slow Controls.
    04:54 ROC 14 reset successful; resuming data-taking.
    05:40 Beam is doing quite well now. We have been running with 100uA with no trips for over an hour now.
    06:05 Starting efficiency run (1636) for f40l kinematics
    06:30 Total statistics for Q2_f40l are roughly 782 events, 4.12% statistical error.
    06:32 Moved HRS-R to new angle (55.36), ready to begin f40r. To minimize target motion the optics run is deferred to the end of the LD2 data.
    06:44 First good run with new kinematics is 1642. 1640 and 1641 are JUNK.
    07:15 - 07:19 q.e. measurement
    07:21 Target IOC crash; secured beam during reboot.
    07:40 Target IOC complete, but the HP heaters may be experiencing a problem. We can only get them to a max of 480 W vice the target of 600W for safe full-current operation. We are coming up to 60 uA to start and will get a target expert.
    07:44 Target looks OK here, trying 75 uA.
    07:50 Beam current at 80 uA. Will operate here for the moment. We have JP on the phone.